OpenStudy (anonymous):

Medal + Fan The height of one solid limestone square pyramid is 24 m. A similar solid limestone square pyramid has a height of 30 m. The volume of the larger pyramid is 16,000 m3. Determine each of the following, showing all your work and reasoning. d) The volume of the smaller pyramid.

OpenStudy (anonymous):

Do you know how to find the volume of a right square pyramid?

OpenStudy (anonymous):

you use the formual V = a^2(h/3) for the large pyramid you have the volume and the height so you would just need to find a. After finding a, you will use the same formula again, however this time you plug in a and they have given you the height so you will solve for the volume.

OpenStudy (anonymous):

4/5^2(24/3)? @Boblovesmath

OpenStudy (anonymous):

But wouldnt it be ^3 since its a 3 dimensional shape?

OpenStudy (anonymous):

I don't know I didn't solve it, but I am sure of the method.

OpenStudy (anonymous):

https://www.google.com/#q=volume+of+a+square+pyramid

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

@SectorZ423 can you figure out the volume?

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

k!

OpenStudy (anonymous):

Get the height of the larger shape and divide it by the smaller one giving you the equation 30/24=

OpenStudy (anonymous):

giving you the number 1.25 and using that to divide the over all volume of the bigger one to get the smaller one so 16,000/1.5=......give me the answer and I will see if its right

OpenStudy (anonymous):

12800

OpenStudy (anonymous):

Okay see thats what my mother got but I was told I was wrong

OpenStudy (anonymous):

the answer you got was wrong its 10,666.67

OpenStudy (anonymous):

Wait no I was wrong

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

Wait so am I right?

OpenStudy (anonymous):

12,800

OpenStudy (anonymous):

OKay Thanks

OpenStudy (anonymous):

yeah no problem I just kinda got switched around with the ratio being 1.5 instead of 1.25